# Visualizing Linear Prediction

### v1

@@ 5 Discussion
There are many other ways to measure the contribution of regressors. Verdinelli and Wasserman (2024) reviewed several, and found that there can be no ‘neutral’ way to do so, and that there is currently no ‘overall best’ way. The displays in this note may seem oversimplified but that makes them easy to interpret. Their main benefit is that they help explainability of the prediction rule as it is given, without depending on how it was derived. It would be possible to extend the displays to some related settings, but that falls outside the scope of this note.
Software availability. The R code and an example script that reproduces all the figures are available at https://wis.kuleuven.be/statdatascience/code/vlp_code_and_script.zip .
Acknowledgments. Mia Hubert and Jakob Raymaekers made helpful comments on the text.
Supplementary Material
@@ Appendix A More figures of the Top Gear data
Figure 2 in the main text was obtained by the commands
fit = lm(1/MPG ~ accel + drive + weight + fuel + 0, data = cars)
main = "Top Gear Data"
grillplot(fit, main = main)
whereas the figures below were generated by
corpred(fit)
car = "Bentley Continental"
grillplot(fit, main = main, casetoshow = car, displaytype = "density",
          xlab = paste0("prediction for ", car))
car = "Kia Rio"
grillplot(fit, main = main, casetoshow = car, staircase = TRUE,
          xlab = paste0("prediction for ", car))
We see the correlations between the four predictors in the first display.
The Bentley Continental is a heavy car with high predicted GPM, whereas the Kia Rio is small and light and requires much less fuel:
@@
This gets picked up automatically in the displays below. Interactions between categorical regressors, or between a numerical and a categorical variable, are visualized in the same way.
@@ Appendix B More figures of the Titanic data
The correlation display below clearly indicates that the passenger’s sex has a large effect on the prediction, followed by passenger class and age. On the other hand, the effect of parch is minuscule. The off-diagonal colors are rather light, indicating smaller correlations between the Titanic predictors than between those of the Top Gear data.
On the next page we see the prediction for passenger 1, whose predicted survival probability is low. Most of the predictors are negative, mainly because the passenger was male and traveled in third class. The only positive predictor comes from his young age. Passenger 2 is a woman traveling in first class, so in spite of the slightly negative age-based predictor her total prediction is high.
@@ Appendix C More figures of the German credit data
Here is the overall grill plot of the German credit data:
@@
newc = list("u.car", 36, 2, 6000, 55, "F", 1)
names(newc) = c("purpose", "months", "rate", "amount", "age", "sex", "nclients")
gp = grillplot(fit, main = main, casetoshow = newc, staircase = TRUE)
@@
> print(gp\$casesummary, row.names = FALSE)
                          predictor    value
@@
            TOTAL LINEAR PREDICTION  2.08699
 TOTAL PREDICTION IN RESPONSE UNITS  0.88963

### v2

@@ 6 Discussion
There are many other ways to measure the contribution of input variables. Verdinelli and Wasserman (2024) reviewed several, and found that there can be no ‘neutral’ way to do so, and that there is currently no ‘overall best’ way. The displays in this note may seem oversimplified but that makes them easy to interpret. Their main benefit is that they help explainability of the prediction rule as it is given, without depending on how it was derived. It would be possible to extend the displays to other settings, but that falls outside the scope of this note.
Software Availability. The R code with an example script that reproduces the figures is available at https://wis.kuleuven.be/statdatascience/code/predsplot_r_code.zip . It will be incorporated in an R package on CRAN.
Acknowledgments. Thanks go to Mia Hubert, Jakob Raymaekers, and the reviewers for helpful suggestions that improved the presentation.
Disclosure Statement. The author reports there are no competing interests to declare.
Supplementary Text
@@ Appendix A More figures of the Top Gear data
We start with the example in the introduction, the fit
$\widehat{hp}=2.466*\mbox{topspeed}-13.13*\mbox{length}+0.063*\mbox{displacement}-206.9$
that was obtained from the standard regression
> fit = lm(hp ~ topspeed + length + displ, data = cars)
> summary(fit)
# Coefficients:
#               Estimate Std. Error t value Pr(>|t|)
# (Intercept) -2.069e+02  2.861e+01  -7.232 4.78e-12 ***
# topspeed     2.466e+00  1.373e-01  17.963  < 2e-16 ***
# length      -1.313e+01  6.187e+00  -2.122   0.0347 *
# displ        6.255e-02  2.808e-03  22.275  < 2e-16 ***
It is a simple example because all of the input variables are numeric. The question was, which of these has the biggest effect on the prediction. Using predsplot() we can easily answer this:
> gp = predsplot(fit, main = "Top Gear data")
> print(gp\$predsummary, row.names = F)
#         prediction term   stdev up/down
#                topspeed  68.380      up
#                  length   5.817    down
#                   displ  91.790      up
#  Total prediction of hp 149.200
We see that the prediction term from disp has the largest standard deviation, followed by topspeed, whereas length only has a tiny effect. The resulting predictions plot visualizes this:
> plot(gp\$p)
We clearly see that the range of the prediction term of length is tiny. This is confirmed by the new display of the correlations and standard deviations of the prediction terms:
> predscor(fit, sort.by.stdev = FALSE)
For Figure 2 in the main text we first ran a standard regression by
> fit = lm(GPM ~ accel + drive + weight + fuel, data = cars)
> summary(fit)
yielding, among other things:
Coefficients:
              Estimate Std. Error t value Pr(>|t|)
(Intercept)  1.681e-02  4.884e-03   3.443 0.000677 ***
accel       -1.356e-03  2.459e-04  -5.516 8.81e-08 ***
driveFront  -2.927e-03  1.619e-03  -1.808 0.071853 .
driveRear   -2.745e-04  1.577e-03  -0.174 0.861992
weight       1.131e-05  1.841e-06   6.143 3.25e-09 ***
fuelPetrol   8.427e-03  1.246e-03   6.761 9.98e-11 ***
The function lm() has encoded the categorical input variable drive with 3 levels by two binary dummy variables driveFront and driveRear, and the categorical variable fuel by the single binary dummy fuelPetrol. From this summary we cannot derive the relative effect of the four input variables accel, drive, weight and fuel.
For the numerical input variables accel and weight there is a known remedy, which is to divide them by their standard deviation and then rerun the fit:
> cars\$st.accel  = cars\$accel/sd(cars\$accel)
> cars\$st.weight = cars\$weight/sd(cars\$weight)
> st.fit = lm(GPM ~ st.accel + drive + st.weight + fuel, data = cars)
> summary(st.fit)
yielding
Coefficients:
              Estimate Std. Error t value Pr(>|t|)
(Intercept)  0.0168141  0.0048841   3.443 0.000677 ***
st.accel    -0.0043289  0.0007848  -5.516 8.81e-08 ***
driveFront  -0.0029267  0.0016189  -1.808 0.071853 .
driveRear   -0.0002745  0.0015774  -0.174 0.861992
st.weight    0.0044897  0.0007308   6.143 3.25e-09 ***
fuelPetrol   0.0084270  0.0012465   6.761 9.98e-11 ***
This changes the coefficients of acceleration and weight, all other coefficients remaining the same. From the absolute values of the new coefficients we see that the effect sizes of acceleration and weight are similar to each other, with that of weight being a bit higher. But we cannot do the same with a categorical variable like drive that is encoded by two dummies. Should we standardize both dummies, and even if we did, how can we combine their effects?
This is where the predictions plot comes in handy. We just run
> main = "Top Gear Data"
> gp = predsplot(fit, main = main)
> plot(gp\$p)
yielding Figure 2, and we get the standard deviations of the prediction terms by
> print(gp\$predsummary, row.names = F)
#          prediction term    stdev up/down
#                    accel 0.004329    down
#                    drive 0.001400
#                   weight 0.004490      up
#                     fuel 0.004104
#  Total prediction of GPM 0.009783
We see that the effect of the input variable drive is the smallest of the four.
The correlations between the four prediction terms are visualized by the line
> predscor(fit)
where we again see that the black square of drive is the smallest.
We can also make a predictions plot for a single case, in order to explain its prediction. To illustrate this we choose two rather extreme cars, the Bentley Continental and the Kia Rio:
> car = "Bentley Continental"
> plot(predsplot(fit, main = main, casetoshow = car, displaytype = "density",
                 xlab = paste0("prediction for ", car))\$p)
> car = "Kia Rio"
> plot(predsplot(fit, main = main, casetoshow = car, staircase = TRUE,
               xlab = paste0("prediction for ", car))\$p)
In the first plot we see that the Bentley gets a very high predicted gallons per mile (GPM). The prediction plot (this time with densities) explains its gas guzzling behavior by the fact that it is heavy (around 2300 kg as seen in the plot), accelerates from standstill to 60 miles per hour in under 5 seconds, runs on petrol, and has all-wheel drive (4WD).
The Kia Rio in the next plot finds itself in the opposite situation, as it is predicted to require very little fuel. The plot (this time in staircase style) explains this by the car’s low weight (around 1200 kg in the plot), slow acceleration, diesel engine, and front wheel drive.
@@
This gets picked up automatically in the displays below. Interactions between categorical input variables, or between a numerical and a categorical variable, are visualized in the same way.
@@ Appendix B More figures of the Titanic data
We now draw the correlation display of the Titanic logistic regression, by the lines
> fit <- glm(y ~ sex + age + sibsp + parch + pclass,family=binomial,data=titanic)
> predscor(fit, sort.by.stdev = FALSE)
It clearly indicates that the passenger’s sex has a large effect on the prediction, followed by passenger class and age. On the other hand, the effect of parch is minuscule. The off-diagonal colors are rather light, indicating smaller correlations between the Titanic prediction terms than between those of the Top Gear data.
The predictions plots for passengers 1 and 2 on the next page are made by
> plot(predsplot(fit, main = main, casetoshow=1)\$p)
> plot(predsplot(fit, main = main, casetoshow=2, staircase = TRUE)\$p)
In the first plot we see that the predicted survival probability of passenger 1 is low. Most of the prediction terms are negative, mainly because the passenger was male and traveled in third class. The only positive prediction term comes from his young age.
Passenger 2 is a woman traveling in first class, so in spite of the slightly negative age-based prediction term her total prediction is relatively high.
@@ Appendix C More figures of the German credit data
Here is the overall predictions plot of the German credit data:
@@
> newc = list("u.car", 36, 2, 6000, 55, "F", 1)
> names(newc) = c("purpose", "months", "rate", "amount", "age", "sex", "nclients")
> gp = predsplot(fit, main = main, casetoshow = newc, staircase = TRUE)
> plot(gp\$p)
@@
> print(gp\$casesummary, row.names = F)
                              prediction term    value
@@
            Total linear prediction of credit  2.08699
 Total prediction of credit in response units  0.88963
Note that non-staircase prediction plots for a single case, like Figure 4, can also be shown with a profile by the argument profile = TRUE :
> plot(predsplot(fit, main = main, casetoshow = 1,
                 displaytype = "density", profile = TRUE)\$p)
The profile is the faint grey broken line that connects the prediction terms of this case. It makes it easier to compare plots of this type made for different cases.